OpenStudy (anonymous):

Evaluate 9 − 3 • 2 + 4, i'm not understanding this, can you please break it down for me in dumbest way possible?

OpenStudy (anonymous):

You do math from left to right. However, (granted that you do not know above multiplication and division) if you see a formula that has multiplication and/or division, you do the multiplication and/or division BEFORE you do the addition and subtraction

OpenStudy (anonymous):

thus in this question, you have a multiplication even though if done in order from left to right, you would obviously start with subtraction. it's just one of the rules of mathematics that you have to get used to.

OpenStudy (anonymous):

the correct order of this question would therefore be: multiplication, subtraction, and addition.

OpenStudy (anonymous):

9 - 3*2 + 4

OpenStudy (anonymous):

9 - 6 + 4

OpenStudy (anonymous):

3 + 4

OpenStudy (anonymous):

7

OpenStudy (anonymous):

So it would be 3*2=6 and then 9-6=3 and then 3+4=7

OpenStudy (anonymous):

yes. sorry i was answering a different question.